OpenStudy (anonymous):

Solving Differential Equation by Separation of variables Find the particular solution y=f(x) to the given differential equation with the initial condition f(1)=2

OpenStudy (anonymous):

dx/dy=xy^3

OpenStudy (anonymous):

dx/x=y^3 dy (y^4)/4=ln x+c y^4=4lnx+C when x=1,y=2 2^4=4ln 1+C 16=0+C C=16 y^4=4lnx+16 ory^4=ln x^4+16